A1

a. Find f(-1), f(1), f(2) and f(3).
b. Find the following limits, if they exist:
i) lim( x->1-, (f(x))
ii) lim( x->1+, (f(x))
iii) lim( x->1, (f(x))
iv) lim( x->2-, (f(x))
v) lim( x->2, (f(x))
c. Is f continuous at x = 1?
d. Find the value of c which makes f continuous at x = 3.
Thankyou!

Found 2 solutions by Aldorozos, MathTherapy:
Answer by Aldorozos(172)

F(-1) = 0
F(1) doesn't exist since by replacing x with 1 we get 0/0 which is undefined
f(2) = 3/4
f(3) = 3 since
i x>-1 = 3
ii 2x-c assuming 1 iii x =1 = f^1 - x
iv = 3/4
v = 3/4
C is not continuous since 0/0 is not defined
d 3*2 - C = 3 Therefore C = 3 (not sure)

Answer by MathTherapy(10552)